. Find a basis for the orthogonal complement of the vector v(1, 3,−2) of the euclidean vector space (R 3 , ·). Argue whether this basis is orthonormal or not.


1
Expert's answer
2022-01-26T16:19:32-0500

Since every vector in the orthogonal complement should be orthogonal to every vector in the given subspace, we need to find the null space of [1, 3, −2]


To find the null space, solve the matrix equation "\begin{bmatrix}
 1 & 3 & -2
\end{bmatrix}""\begin{bmatrix}
 x1 \\
 x2 \\
x3
\end{bmatrix}" = [0].

If we take x2 = t, x3 = s, then x1 = 2s − 3t.

Thus, vector x = "\begin{bmatrix}
 2s - 3t \\
 t \\
s
\end{bmatrix}" = "\begin{bmatrix}
 -3 \\
 1 \\
0
\end{bmatrix}"t + "\begin{bmatrix}
 2 \\
 0 \\
1
\end{bmatrix}"s.

This is the null space.

The basis for the null space is "\begin{Bmatrix}
 \begin{bmatrix}
 -3 \\
 1 \\
0
\end{bmatrix},
\begin{bmatrix}
 2 \\
 0 \\
1
\end{bmatrix}
\end{Bmatrix}".

This is the basis for the orthogonal complement.


Answer: The basis for the orthogonal complement is "\begin{Bmatrix}
 \begin{bmatrix}
 -3 \\
 1 \\
0
\end{bmatrix},
\begin{bmatrix}
 2 \\
 0 \\
1
\end{bmatrix}
\end{Bmatrix}".

To ague whether this basis is orthonormal or not we could find the orthogonal complement of the basis "\begin{Bmatrix}
 \begin{bmatrix}
 -3 \\
 1 \\
0
\end{bmatrix},
\begin{bmatrix}
 2 \\
 0 \\
1
\end{bmatrix}
\end{Bmatrix}" and compare it to the first vector v (1, 3,−2).


Firstly, solve the matrix equation "\begin{bmatrix}
 -3 & 1 & 0 \\
 2 & 0 & 1
\end{bmatrix}""\begin{bmatrix}
 x1 \\
 x2 \\
x3
\end{bmatrix}" = [0].

Write it in matrix form "\begin{bmatrix}
 -3 & 1 & 0 & 0 \\
 2 & 0 & 1 & 0
\end{bmatrix}".


So we have x1 = x1, x2 = 3x1, x3 = -2x1.

So vector x = (1, 3,−2) and equals the vector v (1, 3,−2).

It means that the basis for the orthogonal complement of the vector v was found correctly and is orthonormal.